OpenStudy (anonymous):

. Which of the following points does NOT lie on the graph of the following equation? 5x + 7y = 44 A.(-6, 11) B.(-1, 7) C.(6, 2) D.(13, -3)

OpenStudy (anonymous):

a.(-6, 11)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

@ishaan u r wrong

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

@ishaan94 5(13)+7(-3) =65-21 =44

OpenStudy (anonymous):

ok thnks